OpenStudy (adunb8):

quick question! im trying to find a grade for my class 40% for hw (lowest will be dropped) 10% for quiz 20% for midterm 30% for final 3% for extra credit

OpenStudy (lgbasallote):

that's more than 100%

OpenStudy (lgbasallote):

that totals 103%

OpenStudy (adunb8):

yea but 3% is extra credit so i want to know the total excluding 3%

OpenStudy (lgbasallote):

the total without 3% is 100%

OpenStudy (adunb8):

i have the scores but how do i find each percentage of my scores?

OpenStudy (lgbasallote):

what is your average for your assignment?

OpenStudy (adunb8):

for example i have quiz 1 = 3/3 quiz2 = 24/27 quiz3 = 26/40 quiz4 = 15/15

OpenStudy (adunb8):

so how do i fit it into 10% ?

OpenStudy (lgbasallote):

\[\Large \frac{(\frac 33 \times 100) + (\frac{24}{27} \times 100) + (\frac{26}{40} \times 100) + (\frac{15}{15} \times 100)}{4} \times 0.10\]

OpenStudy (lgbasallote):

it looks complicated...but don't worry...it really is

OpenStudy (adunb8):

oh so i should that for each of them and then add all percentage and i will know right?

OpenStudy (lgbasallote):

like i said...it's complicated....

OpenStudy (lgbasallote):

so just solve that and tell me what you get

OpenStudy (adunb8):

88.47

OpenStudy (lgbasallote):

that's too high...is that just for the quizzes?

OpenStudy (lgbasallote):

your answer for that one should be less than 10

OpenStudy (adunb8):

oops didnt do the x .10 so its 8.847

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

now what about your assignments?

OpenStudy (adunb8):

oh okay thanks! do the same thing and divide by 4 hw assignment and multiply .40 =)

OpenStudy (lgbasallote):

depends on your scores...

OpenStudy (adunb8):

dont i just do the formula you gave me above? but change the division by how many assignments and mulitply by .40 instead of .10?

OpenStudy (lgbasallote):

it's complicated

OpenStudy (adunb8):

hw1 = 232/259 hw2= 36/36 hw3=84/85 hw4=95/100

OpenStudy (lgbasallote):

so what do you think you should do?

OpenStudy (adunb8):

the same way as you did then divide by 4 and muliply by .40?

OpenStudy (lgbasallote):

you have to multiply your scores by 100 first

OpenStudy (lgbasallote):

\[\Large \frac{(\frac{232}{259} \times 100) + (\frac{36}{36} \times 100) + (\frac{84}{85} \times 100) + \frac{95}{100} \times 100)}4 \times 0.40\]

OpenStudy (adunb8):

yes thank you =)))))

OpenStudy (adunb8):

i got it~~~~~

OpenStudy (lgbasallote):

if you're passing grade is 50% then you have a sure pass here

OpenStudy (adunb8):

yes =)

OpenStudy (lgbasallote):

you need a combined score of 6 points for your midterm and finals to pass this subject

OpenStudy (lgbasallote):

note: midterm AND finals